AB INITIO CALCULATIONS OF THE LOW LYING ^{2}Π STATES OF NO

D. W. SCHWENKE, NASA Ames Research Center, Moffett Field, CA 94035-1000.

The NO molecule has too many electrons. The ground electronic state is of ^{2}Π symmetry, and can only form a double bond. The valence electronically excited ^{2}Π states are many, and can only form single bonds or are repulsive. In contrast, the NO^{+} molecule can form a triple bond. Thus there are strongly bound Rydberg states in NO, and these give rise to a myriad of avoided crossings in the Born-Oppenheimer (BO) picture. These crossings are very troublesome, for example the B,C avoided crossing occurs at v = 3 for the C state, and this causes vibrational calculations using the BO picture to be impossible. In this work we show how the BO picture can be de-perturbed in ab initio calculations without resorting to a diabatic transformation starting from the asymptotic region.